Harris owns one share of stock of Fairfax Paint and one share of stock of Litchfield Design. The total value of his holdings is 113 dollars. Both stocks pay annual dividends that are expected to continue forever. The expected return on Fairfax Paint stock is 18.02 percent and its annual dividend is expected to remain at 12.89 dollars forever. What is the next dividend paid by Litchfield Design expected to be if the stock has an annual return of 10.34 percent and dividends are expected to grow by 4.13 percent annually? The next dividend for both firms’ stocks will be paid in one year.
Answer #1

Total value of Harris investment = $113 = stock price of Fairfax Paint and stock price of Litchfield Design

For Fairfax paints

Ke = 18.02%

D = $12.89 forever

using perpetuity, P0 = D/Ke = 12.89/0.1802 = $71.53

So stock price of Litchfield Design = 113-71.53 = $41.47

given for Litchfield Design,

g = 4.13%

Ke = 10.34%

using constant growth model, P0 = D1/(Ke-g)

So, D1 = 41.47*(0.1034-0.0413) = $2.58

next dividend paid by Litchfield Design is expected to be $2.58